OpenStudy (jekdidkdjrjrjjr):

HELP FAN+MEDAL FOR RIGHT ANSWER SEE PIC i need help with c,d,e..... I don't know if c is right so like help pls

OpenStudy (jekdidkdjrjrjjr):

PIC

OpenStudy (jekdidkdjrjrjjr):

HELP ME OMFG

OpenStudy (jekdidkdjrjrjjr):

@dangazzm

OpenStudy (jekdidkdjrjrjjr):

I'll open a new topic for each question so u can get a medal for each one

OpenStudy (anthonyym):

well it doesn't say to use k

OpenStudy (anthonyym):

The problem is that k isn't in the first part of your equation but appears in the second part so it's not equal.

OpenStudy (jekdidkdjrjrjjr):

okay I'm rly bad at this so like

OpenStudy (jekdidkdjrjrjjr):

Like it would be rly nice if u just gave me the answers and I'll give u a medal for each one fair?

OpenStudy (jekdidkdjrjrjjr):

or*

OpenStudy (anthonyym):

You should check this out. It should help: http://www.coolmath.com/algebra/10-complex-numbers/02-factoring-the-sum-of-two-squares-01

OpenStudy (jekdidkdjrjrjjr):

Like I have a lot of questions and a lot of medals to give

OpenStudy (jekdidkdjrjrjjr):

I'm only giving u a medal if u do it sry

OpenStudy (jekdidkdjrjrjjr):

Pls?????

OpenStudy (jekdidkdjrjrjjr):

Like I can open 10 questions and give u 10 medals I'm desperate rn

OpenStudy (dumbcow):

\[a^2 + b^2 = (a+bi)(a-bi)\]

OpenStudy (anthonyym):

Alright. \[x^2 +9\] =(x - 3i)(x - 3i) FOIL that, and it becomes: = x^2+9

OpenStudy (jekdidkdjrjrjjr):

TYTYTY OMFG THATS C RIGHT? I need the others too

OpenStudy (jekdidkdjrjrjjr):

Where do u get the 9 and the 3 from I'm confused

OpenStudy (anthonyym):

In the fundamental theorem of algebra, complex zeros come in pairs. So if you know you have one complex zero, you also have another.

OpenStudy (jekdidkdjrjrjjr):

Domb cows answer looks more legit

OpenStudy (anthonyym):

The 9 and 3 you just randomly picked to prove the identity. You can use a and b if you want.

OpenStudy (jekdidkdjrjrjjr):

Okay ty so D?

OpenStudy (jekdidkdjrjrjjr):

okay chill I have other questions though

OpenStudy (jekdidkdjrjrjjr):

look at da pic

OpenStudy (jekdidkdjrjrjjr):

SO IS ANYONE DOING D OR SHOULD I MAKE A NEW TOPIC ?

OpenStudy (jekdidkdjrjrjjr):

?yo

OpenStudy (anthonyym):

I posted about D about fundamental theorem of algebra previously.

OpenStudy (jekdidkdjrjrjjr):

Wait that wasn't the answer for C?????

OpenStudy (jekdidkdjrjrjjr):

Was the answer you gave me for c or d?

OpenStudy (aishalicious1):

I am not chill I am Aisha and that is some kind of monotone response I wanted to try and wait break it down and type wat u need help with

OpenStudy (jekdidkdjrjrjjr):

I'm so confused rn

OpenStudy (jekdidkdjrjrjjr):

I think I need help with E rn

OpenStudy (jekdidkdjrjrjjr):

And I'm sry I'm just mad bc ive been stuck on dis forever

OpenStudy (aishalicious1):

Type problem e I am not scrolling thru all of these replies to look at the pic

OpenStudy (jekdidkdjrjrjjr):

But like if anyone wants me to make a new topic for more medals it's no prob

OpenStudy (jekdidkdjrjrjjr):

I'll take a pic wait

OpenStudy (jekdidkdjrjrjjr):

PIC I NEED HELP WITH ALL DIS

OpenStudy (aishalicious1):

ALL of it Im not a Math Professor! I'll see wat I can do but Im not a miracle worker

OpenStudy (jekdidkdjrjrjjr):

I don't care if it's a 100% right at this point but like pls try I'm desperate

OpenStudy (aishalicious1):

What is the Fundemental Theory of Algebra? if udk google it and tell me

OpenStudy (jekdidkdjrjrjjr):

Theory idk

OpenStudy (jekdidkdjrjrjjr):

@aishalicious1

OpenStudy (aishalicious1):

Im SO sorry mate had to leave a bunch of friends came around and we went to a party I am sorry hope u dont hate my guts..........